# Redefining Neural Operators in $d+1$ Dimensions

### v2

@@ Abstract
Neural Operators have emerged as powerful tools for learning mappings between function spaces. Among them, the kernel integral operator has been widely validated on universally approximating various operators. Although many advancements following this definition have developed effective modules to better approximate the kernel function defined on the original domain (with $d$ dimensions, $d=1,2,3$...), the unclarified evolving mechanism in the embedding spaces blocks researchers’ view to design neural operators that can fully capture the target system evolution.
Drawing on the Schrödingerisation method in quantum simulations of partial differential equations (PDEs), we elucidate the linear evolution mechanism in neural operators. Based on that, we redefine neural operators on a new $d+1$ dimensional domain. Within this framework, we implement a Schrödingerised Kernel Neural Operator (SKNO) aligning better with the $d+1$ dimensional evolution. In experiments, the $d+1$ dimensional evolving designs in our SKNO consistently outperform other baselines across ten benchmarks of increasing difficulty, ranging from the simple 1D heat equation to the highly nonlinear 3D Rayleigh–Taylor instability. We also validate the resolution-invariance of SKNO on mixing-resolution training and zero-shot super-resolution tasks. In addition, we show the impact of different lifting and recovering operators on the prediction within the redefined NO framework, reflecting the alignment between our model and the underlying $d+1$ dimensional evolution.
@@ 1 Introduction
Neural Operators (NOs) (Kovachki et al. (2023)) parameterize mappings between two function spaces. Specifically, given inputs, they output function values corresponding with adaptable query coordinates in the continuous domain (with $d$ dimensions, $d=1,2,3$...) where functions are originally defined. For operators mapping input to certain output functions defined on a $d$ dimensional domain, there are corresponding $d$ dimensional PDEs describing the governing law between variables of the target system. Once the in/output functions are chosen, the solution operator depends on the governing PDE type. A main challenge for representing different operators is that the corresponding $d$ dimensional PDE types vary across different systems.
Traditional numerical methods (Abdulle (2009); Bangerth (2008)) tackle this by directly converting $d$ dimensional governing PDEs into matrices using techniques like finite element methods and finite difference methods. However, such discretized signal evolving schemes are computationally prohibitive due to their reliance on a finer grid size to improve the solution accuracy (Larson & Bengzon (2013)).
@@
Unlike traditional methods that rely on target system PDEs, common neural networks (NNs) universally learn the mapping between discretized input/output signal pairs on a grid in the $d$ dimensional domain (Hornik et al. (1989); Lagaris et al. (1998)), offering efficiency and adaptability to various governing PDE types. However, their predicted solution is limited by the grid size of signals (Obiols-Sales et al. (2020); Khoo et al. (2021)), suffering from the same problems as numerical methods, i.e., the trade-off between computational costs (training speed, memory allocation, etc.) and accuracy.
@@
Since the problem for common neural networks lies in the size of signal grids, a different philosophy comes: learning operators rather than just matrices, because PDEs govern the evolution of continuous functions rather than discretized signals. A natural extension from matrix to operator is to define a kernel integration to operate on $d$ dimensional functions corresponding to original signals, which has been widely used in modeling neural operators (Li et al. (2021); Liu-Schiaffini et al. (2024); Wu et al. (2024); Li et al. (2020b)).
@@
Specifically, these neural operators empirically introduce an embedding function and evolve it by $d$ dimensional kernel integral operators. Although recent advancements have developed effective modules, such as graph aggregation (Li et al. (2020a; b; 2025)), spectral convolution (Li et al. (2021); Wen et al. (2022)), and attention mechanism (Wu et al. (2024); Xiao et al. (2024)), to better approximate kernel integral operators defined on the $d$ dimensional domain, little discussion for capturing the evolution within the embeddings of input/output functions was given:
Problem 1.
While many works (Kissas et al. (2022); Li et al. (2023b); Rahman et al. (2023)) have evaluated that such operator learning in an embedding space is empirically useful for representing various $d$ dimensional operators, what is the underlying mechanism of introducing an embedding space in PDE / Operator contexts?
Problem 2.
Similar to previous matrix representations, current NO designs are still using discretized matrices to model the evolution within the embeddings. Can we similarly break the spell of unavoidable longer embeddings but heavier computational costs for better approximating?
For problem 1, we try to answer: compared to the matrices corresponding to $d$ dimensional PDEs in common NNs, is there any PDE corresponding to kernel integral operators for the embedding functions in NOs? Here we refer to recent discoveries ((Jin et al. (2024))) in quantum simulation of PDEs: different $d$ dimensional linear PDEs can be converted to their $d+1$ dimensional versions, as a Hamiltonian system in the discrete setting, by using simple warped phase transformations. While common NNs evolve $d$ dimensional signals through matrices, NOs evolve transformed $d+1$ dimensional signals in the embedding space. Both are governed by PDEs with different descriptions of system mechanics. Demonstrations refer to Section 2.1.
For problem 2, we have located the bottleneck for current NOs: In previous common NNs, linear $d$ dimensional PDEs are represented by matrices. For the same system described by $d+1$ dimensional PDEs, neural operators (Li et al. (2021; 2024b); Gupta et al. (2021)) introduce the kernel integral operator to capture the $d$ dimensional evolution, but still utilize matrices for evolving auxiliary embeddings in an Euclidean space. Following this framework, recent advanced methods (Li et al. (2023a); Wu et al. (2024); Li et al. (2025)), especially transformer-based architectures, unavoidably increase the embedding size to better capture the embedding evolution but pay for their computational bill. For example, a compromised way is to use the multi-head to construct a larger block diagonal matrix rather than the complete one which better captures global dependencies during the embedding evolution.
@@
Both problems indicate the same destination: a similar "operator rather than matrix" can be introduced to break the representation bottleneck of current NOs, aligning better with the underlying $d+1$ dimensional evolution. Based on this idea, we firstly demonstrate the $d+1$ dimensional evolving pipe of neural operators, and then redefine neural operators in the $d+1$ dimensional domain (i.e., original $d$ dimensions + one auxiliary dimension). Following the redefined framework, we design the linear block of our model by combining kernel modules in the original domain and along the auxiliary dimension, and then implement the Schrödingerised Kernel Neural Operator (SKNO) within the new framework. Our main contributions in this work are summarized as follows:
@@
  • We redefine the current NO framework where embedding functions are defined on a new $d+1$ dimensional domain according to the Schrödingerisation method (Jin et al. (2024)) in quantum simulation of PDEs, providing a new perspective to understand the evolving mechanism in neural operators, see Sec.2.
@@
  • We design our Schrödingerised Kernel Neural Operator (SKNO) within the $d+1$ dimensional NO framework with linear blocks aligning with the general linear $d+1$ dimensional evolution, utilizing $d+1$ dimensional kernel integral operators with the aid of residuals, see Sec.3.
@@
  • We demonstrate the state-of-the-art performance of our $(d+1)$-dimensional Linear Block and full model across a diverse set of ten benchmarks with varying levels of approximation difficulty. In addition, we conduct model-specific studies, including mixed-training, super-resolution, and ablation experiments, and further explore the impact of different lifting and recovering operators inspired by the observer effect in quantum mechanics.
2 Redefining Neural Operator in $d+1$ Dimensions
2.1 Demonstrations of $d+1$ Dimensional Pipes.
@@ 2 Redefining Neural Operator in $d+1$ Dimensions
To simply clarify the $d+1$ dimensional evolving mechanism related to problem 1, we demonstrate how to transform $d$ dimensional linear PDEs into their $d+1$ dimensional versions with following examples in (Jin et al. (2022; 2023)):
Example 1 (Heat Equation).
@@
Figure 1: A-Schrödingerisation: Classical linear PDEs can be transformed into their $d+1$ dimensional version (here $d=1$ for visualization) by a warped phase transformation, e.g., heat equation $\rightarrow$ phase-space heat equation. And it can be represented by a group of Schrödinger or convection equations after taking the Fourier Transform on $p$ or $x$, see Remark 1. B-Neural Operator: Different $d$ dimensional linear operators can be similarly lifted to their $d+1$ versions, and parameterized by the $d+1$ dimensional Neural Operator which flexibly represents the evolution on both $x$ and $p$.
Step 1: Warped Phase Transformation. Define a transformation from $u(x,t)$ to $v(x,p,t)$ by introducing an auxiliary dimension $p$:
$v(x,p,t=0)=v_{0}(x,p)=e^{-|p|}u_{0},\quad p\in(-\infty,\infty).\qquad\text{($d+1$ dimensional input)}$ (3)
@@
$\partial_{t}v(x,p,t)=-\partial_{p}\partial_{xx}v(x,p,t).\qquad\text{($d+1$ dimensional heat equation)}$ (4)
@@
$u_{1}(x)=\int^{-\infty}_{\infty}\chi(p)v_{1}(x,p)dp,\qquad\text{($d$ dimensional output)}$ (5)
@@
Remark 1 (Fourier Transform of the Phase-space Heat Equation, Fig.1 A).
@@
where $\hat{\tilde{\bm{v}}}(t)$ is derived after taking DFT for $\bm{v}(t)$ on $p$ and $x$, and the same notation $D_{\mu}=diag(\mu_{-M/2},...,\mu_{M/2})$, which denotes a diagonal matrix with $M$ grid points in one dimension, is employed since little confusion will arise. Here, the diagonal coefficient complex matrix $i(D_{\mu}^{2}\otimes D_{\mu})$ is from taking DFTs on the $x$ and $p$ domain for $\partial_{p}\partial_{xx}$ in Equ.3.
@@
Similar process and result on advection equation as Example 1 and Remark 1 with introducing corresponding warped phase transformation $v_{0}=sin(p)u_{0},\ p\in[-\pi,\pi]$. Limited by pages, for more discussion of other PDE types, refer to (Jin et al. (2023)).
According to (Jin et al. (2023); Jin & Liu (2024)), we could generally derive solutions for systems, which are governed by linear PDEs, through the $d+1$ dimensional evolving process as Example 1. In this work, we follow this process to redefine our neural operator framework in $d+1$ dimensions below, shown as Fig.1 B.
@@ 2.2 Operator Learning in $d+1$ Dimensions
Similar to Example 1, we split our framework into three main components based on (Li et al. (2021)) for redefining the $d+1$ dimensional neural operator framework:
@@
To respond to the problem 2 in previous methods, we define the kernel integral operator on the $d+1$ dimensional domain by introducing an auxiliary variable $p$. Define the $d+1$ dimensional kernel integral operator with $\kappa:\mathbb{R}^{d+1}\to\mathbb{R}$:
@@
Note that for non-linear cases, the evolution is usually modeled by iteratively stacking a linear block $\mathcal{L}$ and a non-linear block $\sigma$, which works well in $d$ dimensional NNs (Chen et al. (2018)) and $d+1$ dimensional NOs (He et al. (2024)), shown in Fig.2. With the core component $\mathcal{K}$ in Equ.8, the implementation of $\mathcal{L}$ depends on specific model designs, see Sec.3.1.
@@
Combining the introduced blocks above, now we define our neural operator framework to approximate $\mathcal{G}$ in Sec.2.2 as follows:
Definition 1 (Neural Operator in $d+1$ dimensions).
Assume the $d$ dimensional solution operator $\mathcal{G}$ has at least one corresponding operator in $d+1$ dimensions, define the framework below:
$\mathcal{G}[\bm{a}]$ $=(\mathcal{Q}\circ\mathcal{M}\circ\mathcal{P})[\bm{a}]$
$\approx(\mathcal{Q}\circ\left(\sigma_{L-1}\circ\mathcal{L}_{L-1})\circ\cdots\circ(\sigma_{0}\circ\mathcal{L}_{0}\right)\circ\mathcal{P})[\bm{a}],$ (10)
where $v_{l+1}=(\sigma_{l}\circ\mathcal{L}_{l})[v_{l}],\,l=0,1,...,L-1$. Here, each $\mathcal{L}_{l}$ denotes a general learnable $d+1$ dimensional linear operator, and $\sigma_{l}$ can be a non-linear activation function or a shallow MLP ($p$-dim-evolving operator + activation + $p$-dim-evolving operator).
@@ 3 Schrödingerised Kernel Neural Operator
In this section, we show implementation details of our proposed Schrödingerised Kernel Neural Operator (SKNO) within the redefined $d+1$ dimensional framework, which is informed by the Schrödingerisation process in Sec.2. Sec.3.1 introduces the architecture and implementation of the SKNO, including the modules from preparation lifting to recovery measuring. Different from Sec.3.1 which mainly models the evolution in $D_{x}$, Sec.3.2 focuses on how SKNO better captures the evolution along the auxiliary dimension for its alignment with general linear PDEs.

### v3

@@ Abstract
Neural Operators (NOs) have emerged as powerful tools for learning mappings between function spaces. Among them, the kernel integral operator has been widely used in universally approximating architectures. Following the original formulation, most advancements focus on designing better parameterizations for the kernel over the original physical domain (with $d$ spatial dimensions, $d\in{1,2,3,\ldots}$). In contrast, embedding evolution remains largely unexplored, which often drives models toward brute-force embedding lengthening to improve approximation, but at the cost of substantially increased computation.
In this paper, we introduce an auxiliary dimension that explicitly models embedding evolution in operator form, thereby redefining the NO framework in $d+1$ dimensions (the original $d$ dimensions plus one auxiliary dimension). Under this formulation, we develop a Schrödingerised Kernel Neural Operator (SKNO), which leverages Fourier-based operators to model the $d+1$ dimensional evolution. Across more than ten increasingly challenging benchmarks, ranging from the 1D heat equation to the highly nonlinear 3D Rayleigh–Taylor instability, SKNO consistently outperforms other baselines. We further validate its resolution invariance under mixed-resolution training and super-resolution inference, and evaluate zero-shot generalization to unseen temporal regimes. In addition, we present a broader set of design choices for the lifting and recovery operators, demonstrating their impact on SKNO’s predictive performance.
Machine Learning, ICML
@@ 1 Introduction
Neural Operators (NOs) (Li et al., 2021; Kissas et al., 2022; Kovachki et al., 2023; Wu et al., 2023; Li et al., 2023b; Rahman et al., 2023; Wu et al., 2024b) parameterize mappings between two function spaces. Specifically, given inputs, they output function values corresponding with adaptable query coordinates in the continuous domain (with $d$ dimensions, $d=1,2,3$...) where functions are originally defined. For operators mapping input to certain output functions defined on a $d$ dimensional domain, there are corresponding $d$ dimensional PDEs describing the governing law between variables of the target system. Once the in/output functions are chosen, the solution operator depends on the governing PDE type. A main challenge for representing different operators is that the corresponding $d$ dimensional PDE types vary across different systems.
@@
Traditional numerical methods (Bangerth, 2008; Abdulle, 2009) tackle this by explicitly converting $d$ dimensional governing PDEs into matrices using techniques like finite element methods and finite difference methods. However, such discretized signal evolving schemes are computationally prohibitive when improving the solution accuracy on a finer grid size (Larson and Bengzon, 2013). Unlike traditional methods relying on target system PDEs, common neural networks (NNs) universally learn the mapping between discretized input/output signal pairs on a grid in the $d$ dimensional domain (Hornik et al., 1989; Lagaris et al., 1998), offering efficiency and adaptability to various governing PDE types. However, their predicted solution is limited by the grid size of signals (Obiols-Sales et al., 2020; Khoo et al., 2021), suffering from similar problems as numerical methods, i.e., the trade-off between computational costs (training speed, memory allocation, etc.) and accurate predictions on finer grids.
@@
Since the problem for common neural networks lies in the size of signal grids, neural operators come: learning operators rather than just matrices, because PDEs with certain conditions evolve continuous functions rather than discretized signals. A natural extension from matrix to operator is to define a kernel integration to operate on $d$ dimensional functions fitted to the original signals, which has been widely used in modeling neural operators (Li et al., 2021; Liu-Schiaffini et al., 2024; Wu et al., 2024a; Li et al., 2020b). As a new paradigm for solving PDEs, prior studies (Li et al., 2020a, 2025) have reported that such NOs can achieve errors that are at least an order of magnitude smaller than the common NN.
@@
Specifically, these neural operators empirically introduce a function whose value is an embedding vector, and evolve it by $d$ dimensional kernel integral operators. Keeping activations the same, recent advancements have developed effective modules, such as graph aggregation (Li et al., 2020a, b, 2025), spectral convolution (Li et al., 2021; Wen et al., 2022), and attention mechanism (Wu et al., 2024a; Xiao et al., 2024), to better approximate the linear kernel integral operators defined on the $d$ dimensional domain. However, they have little discussion on modeling the evolution of embeddings from input/output functions: Like the common NN, they utilize matrix multiplication directly on embeddings to evolve them in an Euclidean space by default.
For example, advanced methods (Li et al., 2023a; Wu et al., 2024a; Li et al., 2025), especially transformer-based architectures, often increase the embedding size to better capture embedding evolution, but this incurs a substantial quadratic computational cost. A common compromise is to use multi-heads that effectively form a larger block-diagonal interaction (head-wise mixing) rather than a dense, fully coupled matrix, which would more faithfully capture global dependencies during embedding evolution. However, this head-wise factorization restricts cross-head communication, so global coupling must be recovered indirectly through depth and repeated mixing, further increasing compute.
In summary, current NOs still rely on discretized matrices to model how embeddings evolve. Analogous to the previous success on operator learning, can we utilize operator designs on embeddings to reduce the computational overhead of brute-force embedding lengthening, while preserving (or even improving) approximation quality?
To address this, we depart from the conventional formulation that evolves $d$ dimensional embedding functions over the physical domain. Instead, we explicitly model the evolution as happening simultaneously in (i) the $d$ dimensional physical domain; and (ii) an additional axis representing embedding evolution. Concretely, we introduce an auxiliary dimension and lift the embedding function to a $(d+1)$ dimensional domain, which is inspired by the Schrödingerisation method in quantum simulation of PDEs (Jin et al., 2024). A detailed demonstration is provided in Sec. 2.1.
@@
After clarifying the $d+1$ dimensional evolving pipe of neural operators, we correspondingly redefine neural operators in the $d+1$ dimensional domain (i.e., original $d$ dimensions plus one auxiliary dimension). Following this framework, we implement the Schrödingerised Kernel Neural Operator (SKNO) by combining Fourier-based kernel modules in the original domain and along the auxiliary dimension. Our main contributions in this work are summarized as follows:
@@
  • We redefine the current NO framework where embedding functions are defined on a new $d+1$ dimensional domain, providing a new perspective to model the embedding evolution in neural operators, see Sec.2.
@@
  • We design our Schrödingerised Kernel Neural Operator (SKNO) within the $d+1$ dimensional NO framework with Fourier-based blocks, utilizing $d+1$ dimensional kernel integral operators with the aid of residuals, see Sec.3.
@@
  • We demonstrate the state-of-the-art performance of our SKNO with the $(d+1)$-dimensional kernel integration implementation across a diverse set of more than ten benchmarks with varying levels of approximation difficulty. In addition, we conduct model-specific studies, including mixed-training, super-resolution, zero-shot generalization to unseen temporal regimes, and ablation experiments, and further explore the impact of different lifting and recovering operator choices we provide.
2 Neural Operator in $d+1$ Dimensions
2.1 Demonstrations of $d+1$ Dimensional Evolution
@@
To clarify the $d+1$ dimensional evolving mechanism in Sec.1, we demonstrate how the $d$ dimensional linear PDEs are transformed into their $d+1$ dimensional versions after introducing an auxiliary dimension $p$:
Example 1 (Heat Equation (Jin et al., 2022, 2023)).
@@
Step 1: Warped Phase Transformation. Define a transformation from $u(x,t)$ to $v(x,p,t)$ by introducing an auxiliary dimension $p\in(-\infty,\infty)$:
$v(x,p,t=0)=v_{0}(x,p)=e^{-|p|}u_{0}.$ (3)
@@
$\partial_{t}v(x,p,t)=-\partial_{p}\partial_{xx}v(x,p,t).$ (4)
@@
$u_{1}(x)=\int^{-\infty}_{\infty}\chi(p)v_{1}(x,p)dp,$ (5)
@@
Figure 1: A-Schrödingerisation: Classical linear PDEs can be transformed into their $d+1$ dimensional version (here $d=1$ for visualization) by a warped phase transformation, e.g., heat equation $\rightarrow$ phase-space heat equation. And it can be represented by a group of Schrödinger / convection equations after taking the Fourier Transform on $p$ / $x$. B-Neural Operator: Different $d$ dimensional linear operators can be similarly lifted to their $d+1$ versions, and parameterized by the $d+1$ dimensional Neural Operator which flexibly represents the evolution on both $x$ and $p$.
@@
where $\hat{\tilde{\bm{v}}}(t)$ is derived after taking DFT for $\bm{v}(t)$ on $p$ and $x$, and the same notation $D_{\mu}=diag(\mu_{-N_{p}/2},...,\mu_{N_{p}/2})$, which denotes a diagonal matrix with $N_{p}$ grid points in one dimension, is employed since little confusion will arise. Here, the diagonal coefficient complex matrix $i(D_{\mu}^{2}\otimes D_{\mu})$ is from taking DFTs on the $x$ and $p$ domain for $\partial_{p}\partial_{xx}$ in Equ.3.
@@
Similar process and result on advection equation as Example 1 with introducing corresponding warped phase transformation $v_{0}=sin(p)u_{0},\ p\in[-\pi,\pi]$. Limited by pages, for more discussion of other PDE types, refer to (Jin et al., 2023).
According to (Jin et al., 2023; Jin and Liu, 2024), the warped phase functions and the recovery weight functions along $p$ are chosen based on systems governed by different linear PDEs. Generally, we introduce the auxiliary dimension $p$ with a parameterized setting: explicitly redefining the neural operator embedding functions and their evolution in a $d+1$ dimensional formulation below, as illustrated in Fig. 1 B.
@@ 2.2 Operator Learning in $d+1$ Dimensions
Similar to Example 1, we split our framework into three main components based on (Li et al., 2021) for redefining the $d+1$ dimensional neural operator framework:
@@
We define the kernel integral operator on the $d+1$ dimensional domain by introducing an auxiliary variable $p$. Specifically, given a kernel $\kappa:\mathbb{R}^{2(d+1)}\to\mathbb{R}$, we define:
@@
To assemble a neural operator capable of approximating nonlinear operators, we model the (non-linear) evolution using a standard and well-established structure: iterative stacking of linear operators $\mathcal{L}$ and nonlinear activations $\sigma$, as shown in Fig. 2. When the linear block admits a well-defined operator parameterization, the overall nonlinear approximation can be justified via universal approximation theory (Kovachki et al., 2021). Note that, the concrete implementation of $\mathcal{L}$, whose core component is $\mathcal{K}$ in Eq. 8, depends on the specific model design; see Sec. 3.1.
@@
Given all the definitions introduced above, now we assemble our neural operator framework to approximate $\mathcal{G}$ in the problem setup as follows:
$\mathcal{G}[\bm{a}]$ $=(\mathcal{Q}\circ\mathcal{M}\circ\mathcal{P})[{a}]$
$\approx(\mathcal{Q}\circ\left(\sigma_{L-1}\circ\mathcal{L}_{L-1})\circ\cdots\circ(\sigma_{0}\circ\mathcal{L}_{0}\right)\circ\mathcal{P})[{a}],$ (10)
where $v_{l+1}=(\sigma_{l}\circ\mathcal{L}_{l})[v_{l}],\,l=0,1,...,L-1$. Here, each $\mathcal{L}_{l}$ denotes a general learnable $d+1$ dimensional linear operator, and $\sigma_{l}$ can be a non-linear activation function or a shallow MLP as a non-linear operator ($p$-dim-evolving operator + activation + $p$-dim-evolving operator).
@@ 3 Schrödingerised Kernel Neural Operator
In this section, we show implementation details of our proposed Schrödingerised Kernel Neural Operator (SKNO) within the redefined $d+1$ dimensional framework, which is informed by the Schrödingerisation process in Sec.2. Sec.3.1 introduces the architecture and implementation of the SKNO. Different from Sec.3.1 which mainly follows the evolving in $D_{x}$, Sec.3.2 focuses on how SKNO captures the evolution along the auxiliary dimension.